# A search for Galactic post-AGB stars in Gaia DR3††thanks: Tables A6 and A7 are only available in electronic form at the CDS via anonymous ftp to cdsarc.u-strasbg.fr (130.79.128.5) or via http://cdsweb.u-strasbg.fr/cgi-bin/qcat?J/A+A

### v1

@@ 3 Galactic post-AGB stars identification
The luminosity of a star is a very useful property to distinguish between post-AGB stars and other stellar objects with similar colours, that are located out of the Main Sequence (MS) of the HR diagram, as it occurs with YSOs. To narrow the luminosity range that corresponds to post-AGB stars, we resorted to different evolutionary models for hydrogen-burning post-AGB stars. Concretely, those classical ones by Vassiliadis & Wood (1993) (for masses between 1 and 5 ), and by Bloecker (1995) (for masses between 1 and 7 ), and more recently by Miller Bertolami (2016), which includes different metallicities (for masses between 0.8 and 4 ). Although the range of masses of the progenitor stars is different in each of the models, it should be noted that the evolution of stars with masses greater than 4 is very fast, so rather poor statistics will be expected for objects of these masses and higher. Their expected number should also be very small given the initial mass function. We found a reasonable agreement (see Figure 3) among the models that the range of luminosities that correspond to this evolutionary phase should be .
@@
Such luminosity range for post-AGB stars meets the criteria used by Kamath et al. (2015), who consider that the luminosity range for post-AGB stars is between 2500 and 35,000 . According to these authors, objects above such upper limit may be supergiants or hypergiants, high-mass stars that quickly initiate helium core fusion after they have exhausted their hydrogen, and continue fusing heavier elements after helium exhaustion until they develop an iron core, at which point the core collapses to produce a Type II supernova. In contrast, objects below the lower limit may be post-RGB stars, YSOs, or eventually, other evolved stars such as Horizontal Branch (HB) stars.
@@
The main problem with classifying high luminosity stars as either post-AGB evolved objects or high luminosity massive objects is that both types of objects share many observational features: similar optical spectra, unstable and extended atmospheres, expanding gas-dust envelopes, high IR excesses, and similar IRAS colours (Garcia-Lario et al. 1997). This matter is dealt with in detail by Klochkova & Chentsov (2018), who argue the need to determine and compare various parameters: position in the Galaxy, luminosity, wind parameters, SED, and chemical composition to allow for accurate classification. For this reason, the few objects that we found up the limit will be discussed individually (see Sect. 5).
@@
A different problem occurs with objects below the minimum luminosity indicated by post-AGB evolution models, those with . As we mentioned before, in Kamath et al. (2015) post-AGB candidates in the Magellanic Clouds located in such range were tentatively identified as post-RGB stars. According to these authors, these stars are most likely the result of binary interaction in which their evolution towards the AGB is interrupted, but only in some cases their binary nature has been confirmed, and it could also be the case that some of them are the result of a merging process. It is beyond the scope of this paper to analyse such objects and, for this reason, we would refer to them as unconfirmed post-AGB candidates.
@@
Thus, to classify our objects, as post-AGB stars or another type of stellar objects, it is necessary to calculate accurate luminosities and also to have a reliable estimation of their temperatures.
@@ 3.1 Interstellar reddening
To explore the best possible determination of the interstellar extinction values for our objects, we used Gaia DR3 coordinates and distances and searched the bibliography for the corresponding extinction values. After analysing and contrasting data from different catalogues and dust maps, we decided to use the extinction values from Stassun (2019). Concretely, we obtained values from this catalogue by performing a cross-match between our objects sample and the ’TESS Input Catalog v8.0’ (Stassun 2019), using the Topcat tool (Taylor 2005). The Stassun catalogue contains extinctions from dust maps for 146 objects of our sample. We always used distance-dependent extinction values when they were available. Extinction errors are provided for 116 objects, and around 80% of them are below magnitudes. Interstellar extinction values are listed in Table LABEL:table:candidates.
@@ 3.2 Effective temperatures
We searched the Simbad database for temperature values and references for each object in our sample of 146 candidates and, as a result, we discarded 28 objects that were too hot or with a dubious identification of the central source in the literature. In consequence, we ended up with a sample of 118 objects, that we will consider our final sample. Details are included in Table LABEL:table:candidates. Regarding temperature values, very different situations were found. For approximately 67% of objects, there are precise temperature determinations, which come from spectral analysis. This is the case of 11 objects in common with the work of Kamath et al. (2022), or those in common with Corporaal et al. (2023) or Mello et al. (2012) (see references in Tables A.2, A.3, A.4 and A.5).
@@
For several cases, Teffs come from spectral types derived from medium-resolution spectra, as in the Suárez et al. (2006) catalogue. For others, only Simbad spectral types are available, some of them covering a quite wide range of subtypes or even types. We have also noticed that for some objects, the spectral classifications in the MK system, generally old, are quite discrepant with the spectroscopic temperatures obtained in more recent publications. For instance, the star BD+48 1220 is assigned a spectral type A4Ia (8550 K) in Simbad, from reference Hardorp et al. (1965), while Ting et al. (2019) reports a value of 6389 K from APOGEE spectra analysis. This leads us to deduce that, at least for some cases, the effective temperatures obtained from spectral types may have inaccurate values.
@@
Following the precision of literature values, temperatures from spectral analysis have been prioritised over temperatures obtained by spectral classification in MK types, which in turn have been prioritised over average temperatures obtained directly from the spectral type in the Simbad database. Tables A.2, A.3, A.4, and A.5 list the effective temperature adopted for each object together with a reference and a flag indicating its origin.
@@
3.3 Luminosity and total extinction from SED fitting
@@
To estimate the luminosity of a star, its bolometric flux or magnitude, distance, and interstellar extinction values are needed. A simple approach consists of obtaining the stellar photospheric magnitude V and then applying the bolometric correction to derive the bolometric magnitude. Alternatively, stellar photometry in several bands can be used to build the SED, and then, by fitting it to a certain model, the stellar temperature and luminosity can be predicted. This simple approach is not possible in most cases because of the presence of dust in the interstellar medium, which reddens the spectral distribution and converts the determination of parameters by fitting with a model into a degenerate problem between temperature and extinction.
@@
We will assume as valid the temperature values obtained from the literature with the methodology explained in the previous section. We then used a procedure that allowed us to obtain the luminosity by fitting the SED, introducing the total extinction necessary to obtain the already known temperature value within a range of 250 degrees, which can be considered an acceptable value for the errors of the temperatures assigned from the literature to each object.
@@
The VOSA software is the spanish Virtual Observatory tool designed, among other uses, to estimate effective temperature (Teff), gravity and luminosity based on stellar photometry. The user provides the coordinates of the source, the source distance and its uncertainties and the system looks for observed flux/magnitudes (and their errors) by querying several photometric catalogues accessible through VO services to achieve a wavelength coverage of the data to be analysed as wide as possible. It is possible then to choose among different stellar models to perform the fitting. Concretely, we choose Kurucz models (Castelli & Kurucz, 2003) because they are well fitted for our range of temperatures and the evolutionary stage of post-AGB stars. The VOSA software then performs the absolute flux calibration of the observational data, using the information regarding the available filters (zero points, transmission curves, etc.).
@@
Next, it determines the synthetic photometry for the models with physical parameters in the range selected by the user (in our case, log[g] values between 0 and 5, and metallicity between -4 and 0.5). Dust extinction is also an input to the system, it is provided with no uncertainty together with the selection of an appropriate extinction law. Concretely, VOSA make use of the extinction law by Fitzpatrick (1999) improved by Indebetouw et al. (2005) in the infrared. Next, the best fitting model is provided by VOSA, together with the derivation of the corresponding stellar parameters: Teff and luminosity (as well as a value for log[g], metallicity and overabundances of -elements with respect to iron).
@@
In the problem that interests us, we carried out an iterative procedure which consisted of providing an input test value of the extinction (starting with a value close to the interstellar extinction value from 3D maps) and checking which temperature value was obtained in the SED fitting. We then modified the extinction value in steps of 0.05 magnitudes until a temperature value closer to the literature one within the uncertainty of 250 K mentioned before was obtained. Note that total extinction values obtained from the SED fit can be compared with those from the extinction maps to derive the contribution of the circumstellar component.
@@
Post-AGB candidates in the Torun catalogue were identified as objects with infrared excesses due to the presence of a dust envelope or a disc. This means that those infrared excesses should be accounted for when fitting the SEDs. In VOSA the excesses are detected by iteratively calculating (adding a new data point from the SED at a time) in the mid-infrared (wavelengths redder than 2.5 ) the parameter as defined in Lada et al. (2006). The theoretical spectral models used by VOSA are based on stellar atmospheres and as a result for the calculation of the errors of the fit, the tool only considers those data points of the SED that correspond to bluer wavelengths than the one where the excess has been flagged.
@@
The VOSA service can be used to build SEDs by querying a large variety of photometric surveys available on the platform. In general, we found fluxes for our objects from several catalogues, covering ultraviolet, visible, and near-infrared wavelengths. The most common flux sources used are: 2MASS, DENIS, IRAS, AKARI, WISE, Tycho, Paunzen, UBV, Gaia DR3, Gaia XPy, Pan-Starrs, and GALEX. The flux ranges are shown in Appendix B444https://zenodo.org/uploads/11569760 figures and the individual fluxes for each object are listed in Table A.6 and Table A.7, only available at CDS. The SEDs we provide have been carefully analysed to check for bad data points and then as mentioned before, fitted to Kurucz models (Castelli & Kurucz 2003). We found these models very suited for our sample, as they cover a large temperature range, from 3500 K to 50,000 K.
@@
This procedure has allowed us to derive coherent temperature/total extinction pairs and the luminosity for each of the 118 stars in our final working sample. The uncertainty values in luminosity were estimated by VOSA tool using the uncertainties in the photometry and taking into account the distance uncertainties (lower and upper limits) that we provided as input. The VOSA software provides uncertainties in luminosities, in general, below 10%, which can be considered as a lower limit. As mentioned, VOSA does not support the use of errors in the extinction values to compute the SED fitting. More information about the fitting procedure is provided in Bayo et al. (2008) and in VOSA documentation 555http://svo2.cab.inta-csic.es/theory/vosa.
@@
This procedure, together with accurate distances from Gaia parallaxes, allowed us to obtain values for the luminosities and the total extinctions in agreement, for instance, with those given by Kamath et al. (2022) as it is illustrated in Table 1. However, the fitting procedure used to estimate the luminosity and temperature values has its own limitations, as it does not take into account the uncertainty of the total extinction values (affecting luminosity uncertainties), and it also depends on the fitting models.
@@
The VOSA tool fitting of Kurucz models allows also to obtain tentative values of log[g] (which covers a range between 0 and 5) and metallicity (which covers a range between -4 and 0.5) for our sample stars, although more reliable values for such parameters can be obtained from spectroscopy when available. In Appendix B we provide the fitted SEDs for all the 118 objects in our final sample, including such parameters. Luminosity values together with lower and upper uncertainties, with the limitations explained before, are presented in Tables 3, 4, 5 and 6.
@@
Following Kamath et al. (2015), we analysed the shape of the SEDs and provided a classification in three different types (stellar, shell, or disc). This can give us some additional clues about the possible incidence of binarity in our sample. According to these authors, disc-type SEDs are related to binarity. We classified 30 SEDs as disk-type. The SED morphological classification for our objects is shown in Tables 3, 4, 5 and 6. We opted to locate the disc-type SEDs in an HR diagram along with the rest of the objects as their identification as binaries is tentative and not, in general, confirmed.
@@
Figure 4 depicts the temperature distribution for our final sample objects. Most stars (84%) have values below 10 000 K, as it is expected for post-AGB stars, while, only three stars exhibit effective temperatures above 20 000 K. These last three are sources with infrared flux excesses which have already started to ionise their envelopes, on their way to the PN phase.
@@
If we compare our temperature determinations with those obtained by Oudmaijer et al. (2022) for the 59 objects in common in both samples, we find (Fig. 5) that the temperatures in Oudmaijer et al. (2022) tend to be slightly higher than the ones we obtained. The mean difference being
@@
The temperatures in Oudmaijer et al. (2022) are based on spectral types collected from the Simbad database, which have very different origins and qualities, which might explain the discrepancies.
@@
Regarding the stellar luminosity, Fig. 6 shows that most of the candidate objects have luminosities between , a region that includes the main luminosity range expected for post-AGB stars, but the histogram includes a wide zone of underluminous objects as well.
@@
Figure 7 shows the comparison between our luminosity values and those presented in Oudmaijer et al. (2022) for the 69 stars in common. Note that in Oudmaijer et al. (2022), not all objects with a luminosity value have a temperature value. The luminosity values from Oudmaijer et al. (2022) are very similar to those obtained in this work, the mean difference being
@@
Oudmaijer et al. (2022) determined their luminosities through the dereddened integrated fluxes obtained from Vickers et al. (2015) and multiplying by the square of the distances from Gaia DR3 parallaxes. The authors indicate that the errors in the fluxes are of the order of 20%, which could explain some differences. It is also important to note that Vickers et al. obtained their integrated fluxes assuming default values for the luminosity which implies then that the values in Oudmaijer et al. (2022) are calculated using a rather circular argument.

### v2

@@ 3 Identification of Galactic post-asymptotic giant branch stars
The luminosity of a star is a very useful property for distinguishing between post-AGB stars and other stellar objects with similar colours that are located beyond the main sequence (MS) of the HR diagram, as is the case for YSOs. To narrow the luminosity range that corresponds to post-AGB stars, we resorted to different evolutionary models for hydrogen-burning post-AGB stars: The classical models by Vassiliadis & Wood (1993) (for masses between 1 and 5 ) and by Bloecker (1995) (for masses between 1 and 7 ), and more recently, the model by Miller Bertolami (2016), which includes different metallicities (for masses between 0.8 and 4 ). Although the range of masses of the progenitor stars is different in each of the models, it should be noted that the evolution of stars with masses greater than 4 is very fast. Rather poor statistics is therefore expected for objects of these and higher masses. Their expected number is probably also very small given the initial mass function. The models (see Figure 3) agree reasonably well on a luminosity range of .
@@
This luminosity range for post-AGB stars meets the criteria used by Kamath et al. (2015), who considered that the luminosity range for post-AGB stars is between 2500 and 35,000 . According to these authors, objects above this upper limit may be supergiants or hypergiants, high-mass stars that quickly initiate helium-core fusion after they have exhausted their hydrogen and that continue to fuse heavier elements after helium exhaustion until they develop an iron core, at which point, the core collapses to produce a Type II supernova. In contrast, objects below the lower limit may be post-RGB stars, YSOs, or they may be other evolved stars such as horizontal branch (HB) stars.
@@
The main problem in classifying high-luminosity stars as either post-AGB evolved objects or high-luminosity massive objects is that the two types of objects share many observational features: the optical spectra are similar, they have unstable and extended atmospheres, their gas-dust envelopes expand, they have a high IR excesses, and their IRAS colours are similar (Garcia-Lario et al. 1997). This matter is dealt with in detail by Klochkova & Chentsov (2018), who argued the need to determine and compare various parameters: the position in the Galaxy, the luminosity, the wind parameters, the SED, and the chemical composition to allow for an accurate classification. For this reason, the few objects that we found up to the limit are discussed individually (see Sect. 5).
@@
Objects below the minimum luminosity indicated by post-AGB evolution models present a different problem. These objects have . As we mentioned before, in Kamath et al. (2015), the post-AGB candidates in the Magellanic Clouds located in this range were tentatively identified as post-RGB stars. According to these authors, these stars are most likely the result of binary interaction in which their evolution towards the AGB is interrupted, but only in some cases was their binary nature confirmed, and some of them might equally be the result of a merging process. It is beyond the scope of this paper to analyse these objects, and for this reason, we refer to them as unconfirmed post-AGB candidates.
@@
Thus, to classify our objects as post-AGB stars or as another type of stellar objects, it is necessary to calculate accurate luminosities and also to have reliable estimates of their temperatures.
@@ 3.1 Interstellar reddening
To explore the best possible determination of the interstellar extinction values for our objects, we used Gaia DR3 coordinates and distances and searched the bibliography for the corresponding extinction values. After analysing and comparing data from different catalogues and dust maps, we decided to use the extinction values from Stassun (2019). We obtained values from this catalogue by performing a cross-match between our objects sample and the TESS555Transiting Exoplanet Survey Satellite Input Catalog v8.0 (Stassun 2019) using the Topcat tool (Taylor 2005). The Stassun catalogue contains extinctions from dust maps for 146 objects of our sample. We always used distance-dependent extinction values when they were available. Extinction errors are provided for 116 objects, and the magnitudes of about 80% of them are below . The interstellar extinction values are listed in Table LABEL:table:candidates.
@@ 3.2 Effective temperatures
We searched the Simbad database for temperature values and references for each object in our sample of 146 candidates. As a result, we discarded 28 objects that were too hot or for which the identification of the central source in the literature was uncertain. In consequence, we ended up with a sample of 118 objects, that we considered our final sample. The details are included in Table LABEL:table:candidates. We found very different temperature values. Approximately 67% of objects have precise temperature determinations that come from spectral analysis. This is the case of 11 objects in common with the work of Kamath et al. (2022), or those in common with Corporaal et al. (2023) or Mello et al. (2012) (see references in Tables A.2, A.3, A.4 and A.5).
@@
For several cases, the T effs come from spectral types derived from medium-resolution spectra, as in the Suárez et al. (2006) catalogue. For other cases, only Simbad spectral types are available, some of which cover a quite wide range of subtypes or even types. We also note that for some objects, the spectral classifications in the MK system, which are generally old, are quite discrepant with the spectroscopic temperatures obtained in more recent publications. For instance, the star BD+48 1220 is assigned spectral type A4Ia (8550 K) in Simbad based on Hardorp et al. (1965), while Ting et al. (2019) reported a value of 6389 K from an APOGEE spectra analysis. This leads us to deduce that at least for some cases, the effective temperatures obtained from spectral types may be inaccurate.
@@
Following the precision of the literature values, the temperatures from spectral analysis were prioritised over temperatures obtained by spectral classification in MK types, which in turn were prioritised over average temperatures obtained directly from the spectral type in the Simbad database. Tables A.2, A.3, A.4, and A.5 list the effective temperature we adopted for each object together with a reference and a flag indicating its origin.
@@
3.3 Luminosity and total extinction from fitting the spectral energy distribution
@@
To estimate the luminosity of a star, its bolometric flux or magnitude, distance, and interstellar extinction values are needed. A simple approach consists of obtaining the stellar photospheric magnitude V and then applying the bolometric correction to derive the bolometric magnitude. Alternatively, stellar photometry in several bands can be used to build the SED, and then, by fitting it to a certain model, the stellar temperature and luminosity can be predicted. This simple approach is not possible in most cases because of the dust in the interstellar medium, which reddens the spectral distribution and converts the determination of parameters by fitting with a model into a degenerate problem between temperature and extinction.
@@
We assumed as valid the temperature values obtained from the literature with the method explained in the previous section. We then used a procedure that allowed us to obtain the luminosity by fitting the SED, introducing the total extinction necessary to obtain the already known temperature value within a range of 250 degrees, which can be considered an acceptable value for the errors of the temperatures assigned from the literature to each object.
@@
The VOSA software is the Spanish Virtual Observatory tool that was designed, among other uses, to estimate effective temperature (Teff), gravity, and luminosity based on stellar photometry. The user provides the coordinates of the source, the source distance, and its uncertainties, and the system searches for observed flux (and their errors) by querying several photometric catalogues accessible through VO services to achieve as wide a wavelength coverage of the data to be analysed as possible. We were then able to choose among different stellar models to perform the fitting. We chose Kurucz models (Castelli & Kurucz, 2003) because they are well fitted for our range of temperatures and the evolutionary stage of post-AGB stars. The VOSA software then performed the absolute flux calibration of the observational data, using the information for the available filters (zero points, transmission curves, etc.).
@@
Next, the software determines the synthetic photometry for the models with physical parameters in the range selected by the user (in our case, log[g] values between 0 and 5 and a metallicity between -4 and 0.5). Dust extinction is also an input to the system. It is provided without uncertainty together with the selection of an appropriate extinction law. The VOSA tool makes use of the extinction law by Fitzpatrick (1999) that was improved by Indebetouw et al. (2005) in the infrared. Next, the best-fitting model is provided by VOSA, together with the derivation of the corresponding stellar parameters: Teff and luminosity (and a value for log[g], the metallicity, and the overabundances of -elements with respect to iron).
@@
In the problem that interests us, we carried out an iterative procedure that consisted of providing an input test value of the extinction (starting with a value close to the interstellar extinction value from 3D maps) and determining the temperature value that was obtained in the SED fitting. We then modified the extinction value in steps of 0.05 magnitude until a temperature value closer to the literature value within the uncertainty of 250 K mentioned before was obtained. The total extinction values obtained from the SED fit can be compared with those from the extinction maps to derive the contribution of the circumstellar component.
@@
Post-AGB candidates in the Torun catalogue were identified as objects with an infrared excess due to the presence of a dust envelope or a disc. This means that this infrared excess should be accounted for when fitting the SEDs. In VOSA, the excess is detected by iteratively calculating (adding a new data point from the SED at a time) in the mid-infrared (wavelengths redder than 2.5 ) the parameter as defined in Lada et al. (2006). The theoretical spectral models used by VOSA are based on stellar atmospheres. As a result, the tool only considers the data points of the SED to calculate the fit errors that correspond to bluer wavelengths than the wavelength in which the excess has been flagged.
@@
The VOSA service can be used to build SEDs by querying a large variety of photometric surveys available on the platform. In general, we found fluxes for our objects from several catalogues that covered the ultraviolet, visible, and near-infrared wavelengths. The most common flux sources we used are 2MASS666Two Micron All-Sky Survey, DENIS777Deep Near Infrared Survey, IRAS, spaci, WISE888Wide-field Infrared Survey Explorer, Tycho, Paunzen, UBV999Ultraviolet-Blue-Visible, Gaia DR3, Gaia XPy, Pan-Starrs, and GALEX. The flux ranges are shown in the figures in Appendix B101010https://zenodo.org/uploads/11569760, and the individual fluxes for each object are listed in Table A.6 and Table A.7 (only available at the CDS). The SEDs we provide were analysed to check for bad data points and were then fitted to Kurucz models (Castelli & Kurucz 2003). We found these models very well suited for our sample because they cover a wide temperature range, from 3500 K to 50,000 K.
@@
This procedure has allowed us to derive coherent pairs of temperature/total extinction and the luminosity for each of the 118 stars in our final working sample. The uncertainty values for the luminosity were estimated by VOSA using the uncertainties in the photometry and taking into account the distance uncertainties (lower and upper limits) that we provided as input. The VOSA software provides uncertainties in luminosity below 10% in general, which can be considered as a lower limit. As mentioned, VOSA does not support the use of errors in the extinction values to compute the SED fitting. More information about the fitting procedure is provided in Bayo et al. (2008) and in the VOSA documentation 111111http://svo2.cab.inta-csic.es/theory/vosa/.
@@
This procedure, together with accurate distances from Gaia parallaxes, allowed us to obtain values for the luminosities and the total extinctions that agree, for instance, with those given by Kamath et al. (2022) as illustrated in Table 1. However, the fitting procedure we used to estimate the luminosity and temperature values has its own limitations because it does not take the uncertainty of the total extinction values into account (which affects the luminosity uncertainties), and it also depends on the fitting models.
@@
The VOSA tool fitting of Kurucz models also allowed us to obtain tentative values of log[g] (which covers a range between 0 and 5) and metallicity (which covers a range between -4 and 0.5) for our sample stars, although more reliable values for these parameters can be obtained from spectroscopy when available. In Appendix B we provide the fitted SEDs for all the 118 objects in our final sample, including these parameters. The luminosity values together with lower and upper uncertainties, with the limitations explained before, are presented in Tables 3, 4, 5 and 6.
@@
Following Kamath et al. (2015), we analysed the shape of the SEDs and provide a classification into three different types (stellar, shell, or disc). This can give us some additional clues about the possible incidence of binarity in our sample. According to these authors, disc-type SEDs are related to binarity. We classified 30 SEDs as disc-type. The SED morphological classification for our objects is shown in Tables 3, 4, 5 and 6. We opted to locate the disc-type SEDs in an HR diagram along with the remaining objects as their identification as binaries is tentative and not confirmed in general.
@@
Figure 4 depicts the temperature distribution for our final sample of objects. Most stars (84%) have values below 10 000 K, as expected for post-AGB stars, and only three stars exhibit effective temperatures above 20 000 K. These last three are sources with infrared flux excess that have already started to ionise their envelopes on their way to the planetary nebula phase.
@@
When we compare our temperature determinations with those obtained by Oudmaijer et al. (2022) for the 59 objects in common in both samples, we find (Fig. 5) that the temperatures in Oudmaijer et al. (2022) tend to be slightly higher than those we obtained. The mean difference is
@@
The temperatures in Oudmaijer et al. (2022) are based on spectral types collected from the Simbad database, which have very different origins and qualities. This might explain the discrepancies.
@@
Fig. 6 shows that the stellar luminosities of most of the candidate objects lie between . This region includes the main luminosity range expected for post-AGB stars, but the histogram includes a wide zone of underluminous objects as well.
@@
Figure 7 shows the comparison between our luminosity values and those presented in Oudmaijer et al. (2022) for the 69 stars in common. Not all objects with a luminosity value in Oudmaijer et al. (2022) have a temperature value. The luminosity values from Oudmaijer et al. (2022) are very similar to those obtained in this work, the mean difference is
@@
Oudmaijer et al. (2022) determined their luminosities through the dereddened integrated fluxes obtained from Vickers et al. (2015) and by multiplying by the square of the distances from Gaia DR3 parallaxes. The authors indicated that the errors in the fluxes are about 20%, which could explain some differences. It is also important to note that Vickers et al. obtained their integrated fluxes assuming default values for the luminosity, which implies then that the values in Oudmaijer et al. (2022) were calculated using a rather circular argument.